OpenStudy (anonymous):

f ( x ) = 6 - 9x^2 I need step by step please

OpenStudy (anonymous):

@Hero

hero (hero):

What are the instructions for this problem?

OpenStudy (anonymous):

@Hero please solve for f(9)

hero (hero):

f(x) = 6 - 9x^2 f(9) = 6 - 9(9)^2 Simplify the RHS

OpenStudy (anonymous):

What's a RHS?

hero (hero):

Simplify the right hand side.

OpenStudy (anonymous):

Okay, what next?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@mathmale please help?

hero (hero):

Have you simplified the expression yet? Show your work.